# Low-Latency Sliding Window Connectivity

### v1

@@ Abstract.
PVLDB Reference Format:
PVLDB, 18(1): XXX-XXX, 2025.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/chaozhang-cs/lswc.
@@ 1. INTRODUCTION
Connectivity queries that check whether vertices belong to the same connected components are fundamental to graph computations (Sahu et al., 2017, 2020). With the prevalence of graph representations in numerous real-world domains (Newman, 2010; Sakr et al., 2021), these queries are critical for a variety of applications, including social networks (Otte and Rousseau, 2002), transport networks (Von Ferber et al., 2009), financial networks (Hodler, 2023), and more.
@@
Figure 1. Running example.
Running example. Figure 1(a) presents a running example of sliding windows with a window size of $5$ timestamps and a slide interval of $1$ timestamp, where each timestamp could represent a time unit such as a minute. It shows a window snapshot $w_{7}$, containing $5$ timestamps, and illustrates the corresponding graph $\mathcal{G}(w_{7})$. When the window moves expired edges (with timestamp $t_{7}$) are deleted from the window and new edges (with timestamp $t_{12}$) are inserted, transitioning from $w_{7}$ to $w_{8}$ and generating $\mathcal{G}(w_{8})$. The query result for vertices $I$ and $G$ in both $\mathcal{G}(w_{7})$ and $\mathcal{G}(w_{8})$ are True, as they belong to the same connected component.
@@
Alternatively, an index-based strategy can be used that involves using data structures designed for Fully Dynamic Connectivity (FDC) (Iyer et al., 2002; Henzinger and King, 1995; Holm et al., 1998; Henzinger and King, 1999; Eppstein et al., 1997; Holm et al., 2001; Thorup, 2000; Henzinger and King, 2001; Wulff-Nilsen, 2013; Kapron et al., 2013; Kejlberg-Rasmussen et al., 2016; Huang et al., 2017; Wulff-Nilsen, 2017; Patrascu and Demaine, 2006; Henzinger and Fredman, 1998; Chen et al., 2022; Thorup, 2000). These techniques maintain spanning trees within windows, with one tree per connected component, supporting query processing, edge insertion, and edge deletion. Connectivity queries for $(u,v)$ can be efficiently answered by checking whether $u$ and $v$ belong to the same spanning tree, which involves simply traversing from each node to its root. New edge insertions can also be processed efficiently by linking spanning trees when the two endpoints of the new edge are not already connected. However, significant performance challenges arise when updating spanning trees due to the deletion of expired edges. Deleting an edge from a spanning tree requires searching for replacement edges (see §3.2) to reconnect the resulting subtrees, which require a full graph traversal in the worst-case scenario. In the running example, Figure 1(b) depicts the spanning tree in $\mathcal{G}(w_{7})$, and the expired edge $(B,D)$ has a replacement edge $(E,F)$.
@@
The main idea of our framework is that a maximum spanning tree is maintained for each connected component in each window, using edge timestamps as weights, allowing for the straightforward deletion of expired edges. For instance, two spanning trees for tracking the connected components in $\mathcal{G}(w_{7})$ are shown in Figures 1(b) and (c). When $w_{7}$ is transitioning to $w_{8}$, deleting the expired edge $(B,D)$ in the spanning tree in Figure 1(b) results in checking the existence of a non-tree edge (dotted in red) that can reconnect vertices $B$ and $D$, e.g., edge $(E,F)$, which is an expensive operation due to required graph traversal. The deletion in the maximum spanning tree in Figure 1(c), on the other hand, does not require such expensive computation or any additional operations besides the deletion itself.
@@
Due to space constraints, all proofs are available in our online technical report (Zhang et al., 2024b).
@@ 2. RELATED WORK
Fully dynamic connectivity. FDC (Iyer et al., 2002; Henzinger and King, 1995; Holm et al., 1998; Henzinger and King, 1999; Eppstein et al., 1997; Holm et al., 2001; Thorup, 2000; Henzinger and King, 2001; Wulff-Nilsen, 2013; Kapron et al., 2013; Kejlberg-Rasmussen et al., 2016; Huang et al., 2017; Wulff-Nilsen, 2017; Patrascu and Demaine, 2006; Henzinger and Fredman, 1998; Chen et al., 2022; Thorup, 2000; Hanauer et al., 2022) involves maintaining connectivity information over graphs undergoing arbitrary updates, including both edge insertions and deletions. Data structures designed for computing FDC support three operations: insert, delete, and query. Typically, spanning trees are maintained internally, with each processed edge classified as either a tree edge or a non-tree edge. Both insertion and deletion operations for either type of edges can be efficiently computed, except for the deletion of a tree edge, which requires a replacement edge search (see §3.2). This search can take $O(|V|+|E|)$ time in the worst case and is a major performance bottleneck of all existing FDC approaches. While FDC data structures can be used to compute sliding window connectivity (as they inherently support insert, delete, and query), the performance bottleneck becomes more severe in this case because every edge inserted into the window will eventually expire and need to be deleted. In this work, we design a framework that leverages the benefits of FDC data structures for query processing while avoiding the expensive replacement edge search. Our experiments show that integrating D-Tree, the most efficient FDC solution in practice, with our framework significantly outperforms vanilla D-Tree (Chen et al., 2022).
@@
Stream processing systems. High throughput and low latency are crucial in stream processing (Toshniwal et al., 2014; Kulkarni et al., 2015; Akidau et al., 2013; Zaharia et al., 2013; Carbone et al., 2015). The continuous stream processing model, used in systems like Apache Flink (Carbone et al., 2015), processes queries in real-time as windows complete. We consider this model for processing sliding window connectivity queries. Although various streaming systems for graph computations exist (Mariappan and Vora, 2019; Sheng et al., 2018; Iyer et al., 2016; Sengupta et al., 2016; Kumar and Huang, 2020; Ediger et al., 2012; Barbieri et al., 2009; Le-Phuoc et al., 2011; Calbimonte et al., 2010; Dell’Aglio et al., 2015; Rost et al., 2021), they lack specific indexing methods for connectivity query processing. We aim to address this gap and efficiently support these fundamental graph computations.
@@ 3.1. Sliding Window Connectivity
Streaming graphs in sliding windows. A sliding window maintains a fixed-size subset of a continuous data stream, denoted as $w$. In this paper, we use a time-based window model with a fixed window size ($\alpha$) and slide intervals ($\beta$), both in time units. This defines a sequence of window snapshots $(w_{1},w_{2},\ldots)$, where each $w_{i}$ has a beginning timestamp $w_{i}.t_{b}$ and an ending timestamp $w_{i}.t_{e}$, containing $\alpha$ time units. For adjacent snapshots $w_{i}$ and $w_{i+1}$, $w_{i+1}.t_{b}=w_{i}.t_{b}+\beta$. Each $w$ contains all edges $e_{i}$ where $w.t_{b}\leq e_{i}.t\leq w.t_{e}$, and the graph in $w$ is denoted as $\mathcal{G}(w)$. Since $\alpha$ is larger than $\beta$ in the context of sliding windows, window snapshots overlap.
@@ 4.3. Inserting New Edges
When inserting a streaming new edge $(u,v)$ into an MST in any window snapshot of a sliding window, and $u$ and $v$ are already connected in the MST, the spanning tree will remain an MST by converting the minimum tree edge [^1] If the minimum edges are not unique, any of them can be selected for the converting. in the cycle containing $(u,v)$ into a non-tree edge and then inserting $(u,v)$ as a tree edge.
@@
Algorithm 1 Sliding Window Connectivity with MSTs
@@ 4.4. Sliding Window Connectivity with MSTs
Query $Q_{c}(u,v)$ can be simply computed by checking whether $u$ and $v$ belong to the same MST. Additionally, for each new edge $(u,v)$, we determine if the new edge should be treated as a tree edge by verifying if $Q_{c}(u,v)=\text{False}$ before inserting the new edge. If not, then the new edge is a non-tree edge.
@@
In the remainder of the paper, we describe how to use FDC data structures to implement the abstract data types of MSTs required in Algorithm 1 in Section 5. Additionally, we discuss optimized implementations in Section 6 using the most advanced data structures from both practical and theoretical perspectives.
@@ 5.1. Computing MSTs
To summarize, any data structure designed to handle FDC can be adapted to compute MST in Algorithm 1. It is important to note that FDC data structures can be directly used to compute sliding window connectivity, as they inherently support the insert, delete, and query operations required for sliding window connectivity. We refer to such direct adoption as vanilla solutions. In contrast, integrating FDC data structures to compute MSTs in Algorithm 1 for computing sliding window connectivity is referred to as MST solutions. The main difference between the vanilla solution and the MST solution is that, in the former case, it is necessary to search for replacement edges when deleting tree edges, a step that the MST solution eliminates. Although the MST solution requires the additional findMinInCycle operation, it is significantly more efficient than the vanilla solution in terms of both throughput and latency, as demonstrated by our experimental evaluation.
@@ 6.5. Complexity Analysis
OMST S-Tree has $O(n)$ worst-case and amortized time complexity for query, insert, and delete due to traversing paths from vertices to their roots. Deletion is also $O(n)$ because it involves updating vertex sizes along these paths. OMST D-Tree and MST D-Tree have the same worst-case time complexity as OMST S-Tree for the same reason. However, OMST D-Tree and MST D-Tree are expected to have better average-case complexity due to applying D-Tree’s Techniques 1 and 2, reducing the path length from vertices to their roots on average (see D-Tree (Chen et al., 2022) for details). OMST LC-Tree can achieve $O(\log{n})$ amortized time for each of query, insert, and delete because all these operations are executed by a few invocations of the access operation, taking $O(\log n)$ amortized time (Sleator and Tarjan, 1981; Tarjan and Werneck, 2010) when using splay trees to store preferred paths.
@@ 7.1. Experimental Setup
The following $8$ approaches are compared in our experimental evaluation: Depth-First Search (DFS), RWC, Vanilla D-Tree (Chen et al., 2022), BIC (Zhang et al., 2024a), MST D-Tree, OMST S-Tree, OMST D-Tree, and OMST LC-Tree. Among these, DFS and RWC are non-indexing approaches, while the others are indexing approaches. The DFS approach applies depth-first search to compute each query in each window. RWC recomputes the connected components in each window and uses them to process queries. Vanilla D-Tree uses D-Tree as the FDC data structure to compute sliding window connectivity. We note that other FDC data structures, e.g., Euler-Tour Tree (Tarjan and Vishkin, 1985) and HDT (Holm et al., 1998, 2001), are excluded from our experimental evaluation due to their suboptimal performance, as demonstrated previously (Zhang et al., 2024a). MST D-Tree (§5.2), OMST S-Tree (§6.2), OMST D-Tree (§6.3), and OMST LC-Tree (§6.4) are the approaches proposed in this paper.
@@ 7.2. Tail Latency and Throughput
Figure 5 shows P95 and P99 latency for both query and window management, while Figure 6 reports throughput. Both experiments use windows with an average size of $3$M edges, slide intervals averaging $150$K edges, and workloads of $10^{5}$ queries.
@@ 7.5. Memory Usage
We present the experiment results on memory consumption in Figure 10: 10(a) details memory usage for each graph with a window size of 3M edges and a slide size of 150K edges; 10(b)-(d) illustrate memory usage with a fixed slide size of 1M edges across various window sizes; 10(e)-(f) depict memory usage for a fixed window size of 80M edges with varying slide sizes. Overall, OMST D-Tree and OMST S-Tree use the least amount of memory among index-based approaches. RWC consumes the least amount of memory because it recomputes connected components for each window, resulting in memory usage linear to the number of vertices in windows. OMST D-Tree, OMST S-Tree, and OMST LC-Tree consume less memory than MST D-Tree, as they store only spanning trees, while MST D-Tree also stores non-tree edges, similar to Vanilla D-Tree. OMST D-Tree and OMST S-Tree have identical memory usage, less than OMST LC-Tree, which requires extra pointers for its access operation. BIC, at least storing all streaming edges in each disjoint chunk that contains all streaming edges in a window, generally uses more memory than OMST-based approaches. Memory usage increases with window size for all but the OMST framework approaches; slide size does not affect memory usage. Although both RWC and the OMST framework approaches (OMST D-Tree, OMST S-Tree, and OMST LC-Tree) have the same space complexity of $O(n)$, where $n$ is the number of vertices in a window, their actual memory usage differs. In the OMST framework, only a parent pointer is maintained for each vertex in spanning trees. When deleting expired edges, if a vertex is no longer connected to any edges in the spanning tree, it is not removed. Deleting such vertices would require maintaining children pointers for each vertex, effectively doubling the storage requirements of spanning trees. As a result, the OMST framework might store more vertices than RWC. However, in scenarios with large window sizes, the memory usage between the OMST framework and RWC becomes similar, as shown in Figures 10(b)-(d). We note that maintaining children pointers for each vertex is straightforward and does not overly complicate the structure.

### v2

@@ 1. INTRODUCTION
Connectivity queries that check whether vertices belong to the same connected components (CCs) are fundamental to graph computations (Sahu et al., 2017, 2020). With the prevalence of graph representations in numerous real-world domains (Newman, 2010; Sakr et al., 2021), these queries are critical for a variety of applications, including social networks (Otte and Rousseau, 2002), transport networks (Von Ferber et al., 2009), financial networks (Hodler, 2023), and more.
@@
Figure 1. Running example of sliding window connectivity.
Example. Figure 1(a) presents a running example of sliding windows with a window size of $5$ timestamps and a slide interval of $1$ timestamp, where each timestamp could represent a time unit such as a minute. It shows window snapshot $w_{7}$ containing edges with timestamps from $t_{7}$ to $t_{11}$ and illustrates the corresponding graph $\mathcal{G}(w_{7})$. When the window is moving from $w_{7}$ to $w_{8}$, expired edges (with timestamp $t_{7}$) will be deleted from $\mathcal{G}(w_{7})$ and new edges (with timestamp $t_{12}$) are inserted, generating $\mathcal{G}(w_{8})$. The query result for vertices $I$ and $G$ in both $\mathcal{G}(w_{7})$ and $\mathcal{G}(w_{8})$ are True, as they belong to the same CC in $\mathcal{G}(w_{7})$ and $\mathcal{G}(w_{8})$, respectively.
@@
Alternatively, an index-based strategy can be used that involves using data structures designed for Fully Dynamic Connectivity (FDC) (Iyer et al., 2002; Henzinger and King, 1995; Holm et al., 1998; Henzinger and King, 1999; Eppstein et al., 1997; Holm et al., 2001; Thorup, 2000; Henzinger and King, 2001; Wulff-Nilsen, 2013; Kapron et al., 2013; Kejlberg-Rasmussen et al., 2016; Huang et al., 2017; Wulff-Nilsen, 2017; Patrascu and Demaine, 2006; Henzinger and Fredman, 1998; Chen et al., 2022; Thorup, 2000). These techniques maintain spanning trees within windows, with one tree per CC, supporting query processing, edge insertion, and edge deletion. Connectivity queries for $(u,v)$ can be efficiently answered by checking whether $u$ and $v$ belong to the same spanning tree, which involves simply traversing from each node to its root. New edge insertions can also be processed efficiently by linking spanning trees when the two endpoints of the new edge are not already connected. However, significant performance challenges arise when updating spanning trees due to the deletion of expired edges. Deleting an edge from a spanning tree requires searching for replacement edges (see §3.2) to reconnect the resulting subtrees, which require a full graph traversal in the worst-case scenario. In the running example, Figure 1(b) depicts the spanning tree in $\mathcal{G}(w_{7})$, and the expired edge $(B,D)$ has a replacement edge $(E,F)$.
@@
The main idea of our framework is that a maximum spanning tree is maintained for each CC in each window, using edge timestamps as weights, allowing for the straightforward deletion of expired edges. For instance, two spanning trees for tracking the CCs in $\mathcal{G}(w_{7})$ are shown in Figures 1(b) and (c). When $w_{7}$ is transitioning to $w_{8}$, deleting the expired edge $(B,D)$ in the spanning tree in Figure 1(b) results in checking the existence of a non-tree edge (dotted in red) that can reconnect vertices $B$ and $D$, e.g., edge $(E,F)$, which is an expensive operation due to required graph traversal. The deletion in the maximum spanning tree in Figure 1(c), on the other hand, does not require such expensive computation or any additional operations besides the deletion itself.
@@ 2. RELATED WORK
Fully dynamic connectivity. FDC (Iyer et al., 2002; Henzinger and King, 1995; Holm et al., 1998; Henzinger and King, 1999; Eppstein et al., 1997; Holm et al., 2001; Thorup, 2000; Henzinger and King, 2001; Wulff-Nilsen, 2013; Kapron et al., 2013; Kejlberg-Rasmussen et al., 2016; Huang et al., 2017; Wulff-Nilsen, 2017; Patrascu and Demaine, 2006; Henzinger and Fredman, 1998; Chen et al., 2022; Thorup, 2000; Hanauer et al., 2022) involves maintaining connectivity information over graphs undergoing both edge insertions and deletions. Data structures designed for computing FDC support three operations: insert, delete, and query. Spanning trees are maintained internally, with each processed edge classified as either a tree edge or a non-tree edge. Both insertion and deletion operations for either type of edges can be efficiently computed, except for the deletion of a tree edge, which requires a replacement edge search (see §3.2). This search can take $O(|V|+|E|)$ time in the worst case and is a major performance bottleneck of all existing FDC approaches. While FDC data structures can be used to compute sliding window connectivity (as they inherently support insert, delete, and query), the performance bottleneck becomes more severe in this case because every edge inserted into the window will eventually expire and need to be deleted. In this work, we design a framework that leverages the benefits of FDC data structures for query processing while avoiding the expensive replacement edge search. Our experiments show that integrating D-Tree, the most efficient FDC solution in practice, with our framework significantly outperforms vanilla D-Tree (Chen et al., 2022).
@@
Stream processing systems. High throughput and low latency are crucial in stream processing (Toshniwal et al., 2014; Kulkarni et al., 2015; Akidau et al., 2013; Zaharia et al., 2013; Carbone et al., 2015). The continuous stream processing model, used in systems like Apache Flink (Carbone et al., 2015), processes queries in real-time as windows complete. We consider this model for processing sliding window connectivity queries. Although various streaming systems for graph computations exist (Mariappan and Vora, 2019; Sheng et al., 2018; Iyer et al., 2016; Sengupta et al., 2016; Kumar and Huang, 2020; Ediger et al., 2012; Barbieri et al., 2009; Le-Phuoc et al., 2011; Calbimonte et al., 2010; Dell’Aglio et al., 2015; Rost et al., 2021), they lack specific indexing methods for connectivity query. We aim to address this gap and efficiently support these fundamental graph computations.
@@ 3.1. Sliding Window Connectivity
Streaming graphs in sliding windows. A sliding window maintains a fixed-size subset of a continuous data stream, denoted as $w$. In this paper, we use a time-based window model with a fixed window size ($\alpha$) and slide intervals ($\beta$), both in time units. This defines a sequence of window snapshots $(w_{1},w_{2},\ldots)$, where each $w_{i}$ has a beginning timestamp $w_{i}.t_{b}$ and an ending timestamp $w_{i}.t_{e}$, containing $\alpha$ time units. For adjacent snapshots $w_{i}$ and $w_{i+1}$, $w_{i+1}.t_{b}=w_{i}.t_{b}+\beta$ and $w_{i+1}.t_{e}=w_{i}.t_{e}+\beta$. Each $w$ contains all edges $e_{i}$ where $w.t_{b}\leq e_{i}.t\leq w.t_{e}$, and the graph in $w$ is denoted as $\mathcal{G}(w)$. Since $\alpha$ is larger than $\beta$ in the context of sliding windows, window snapshots overlap.
@@ 4.3. Inserting New Edges
When inserting a streaming new edge $(u,v)$ into an MST in any window snapshot of a sliding window, and $u$ and $v$ are already connected in the MST, the spanning tree will remain an MST by converting the minimum tree edge [^1] If the minimum edges are not unique, any can be selected for converting. in the cycle containing $(u,v)$ into a non-tree edge and then inserting $(u,v)$ as a tree edge.
@@
Algorithm 1 The MST Framework
@@ 4.4. Sliding Window Connectivity with MSTs
Query $Q_{c}(u,v)$ can be computed by checking whether $u$ and $v$ belong to the same MST. Additionally, for each new edge $(u,v)$, we determine if the new edge should be treated as a tree edge by verifying if $Q_{c}(u,v)=\text{False}$ before inserting the new edge. If not, then the new edge is a non-tree edge.
@@
In the remainder of the paper, we describe how to use FDC data structures to implement the abstract data types of MSTs required in Algorithm 1 in Section 5. Then, we discuss optimized implementations in Section 6 using the most advanced techniques.
@@ 5.1. Computing MSTs
To summarize, any FDC data structure can be adapted to compute MST in Algorithm 1. It is important to note that FDC data structures can also be used to compute sliding window connectivity without adaption, as they inherently support insert, delete, and query required for sliding window connectivity. We refer to such approaches as vanilla solutions. In contrast, integrating FDC data structures to compute MSTs in Algorithm 1 for computing sliding window connectivity is referred to as MST solutions. The main difference between the vanilla solution and the MST solution is that, in the former case, it is necessary to search for replacement edges when deleting tree edges, a step that the MST solution eliminates. Although the MST solution requires the additional findMinInCycle operation, it is significantly more efficient than the vanilla solution in terms of both throughput and latency, as demonstrated by our experimental evaluation.
@@ 6.5. Complexity Analysis
OMST S-Tree has $O(n)$ worst-case and amortized time complexity for query, insert, and delete due to traversing paths from vertices to their roots. Deletion is also $O(n)$ because it involves updating vertex sizes along these paths. OMST D-Tree and MST D-Tree have the same worst-case time complexity as OMST S-Tree for the same reason. However, OMST D-Tree and MST D-Tree are expected to have better average-case complexity due to applying D-Tree’s Techniques 1 and 2, reducing the path length from vertices to their roots on average (see D-Tree (Chen et al., 2022) for details). OMST LC-Tree can achieve $O(\log{n})$ amortized time for each of query, insert, and delete because all these operations are executed by a constant number of invocations of the access operation, taking $O(\log n)$ amortized time (Sleator and Tarjan, 1981; Tarjan and Werneck, 2010) when using splay trees to store preferred paths (see LC-Tree (Sleator and Tarjan, 1981) for details).
@@ 7.1. Experimental Setup
The following $8$ approaches are compared in our experimental evaluation: Depth-First Search (DFS), RWC, Vanilla D-Tree (Chen et al., 2022), BIC (Zhang et al., 2024a), MST D-Tree, OMST S-Tree, OMST D-Tree, and OMST LC-Tree. Among these, DFS and RWC are non-indexing approaches, while the others are indexing approaches. The DFS approach applies depth-first search to compute each query in each window. RWC recomputes the connected components in each window and uses them to process queries. Vanilla D-Tree uses D-Tree as the FDC data structure to compute sliding window connectivity. We note that other FDC data structures, e.g., Euler-Tour Tree (Tarjan and Vishkin, 1985) and HDT (Holm et al., 1998, 2001), are excluded from our experimental evaluation due to their suboptimal performance compared to BIC and Vanilla D-Tree, as demonstrated previously (Zhang et al., 2024a). MST D-Tree (§5.2), OMST S-Tree (§6.2), OMST D-Tree (§6.3), and OMST LC-Tree (§6.4) are the approaches proposed in this paper. Our codebase [^2] Open Source Link: https://github.com/chaozhang-cs/lswc for implementation and experimental evaluation has been made publicly available online.
@@ 7.2. Tail Latency and Throughput
Figure 5 shows P95 and P99 latency for query and window management, while Figure 6 reports throughput. Both experiments use windows with an average size of $3$M edges, slide intervals averaging $150$K edges, and workloads of $10^{5}$ queries.
@@ 7.5. Memory Usage
Memory consumption results are presented in Figure 10: 10(a) details memory usage for each graph with a window size of 3M edges and a slide size of 150K edges; 10(b)-(d) illustrate memory usage with a fixed slide size of 1M edges across various window sizes; 10(e)-(f) depict memory usage for a fixed window size of 80M edges with varying slide sizes. Overall, OMST D-Tree and OMST S-Tree use the least amount of memory among index-based approaches. RWC consumes the least amount of memory because it recomputes connected components for each window, resulting in memory usage linear to the number of vertices in windows. OMST D-Tree, OMST S-Tree, and OMST LC-Tree consume less memory than MST D-Tree, as they store only spanning trees, while MST D-Tree also stores non-tree edges, similar to Vanilla D-Tree. OMST D-Tree and OMST S-Tree have identical memory usage, less than OMST LC-Tree, which requires extra pointers for its access operation. BIC, at least storing all streaming edges in each disjoint chunk that contains all streaming edges in a window, generally uses more memory than OMST-based approaches. Memory usage increases with window size for all but the OMST framework approaches; slide size does not affect memory usage. Although both RWC and the OMST framework approaches (OMST D-Tree, OMST S-Tree, and OMST LC-Tree) have the same space complexity of $O(n)$, where $n$ is the number of vertices in a window, their actual memory usage differs. In the OMST framework, only a parent pointer is maintained for each vertex in spanning trees. When deleting expired edges, if a vertex is no longer connected to any edges in the spanning tree, it is not removed. Deleting such vertices would require maintaining child pointers for each vertex, effectively doubling the storage requirements. As a result, the OMST framework might store more vertices than RWC. However, in scenarios with large window sizes, the memory usage between the OMST framework and RWC becomes similar, as shown in Figures 10(b)-(d). We note that maintaining child pointers for each vertex is straightforward and does not overly complicate the structure.